A machine has a book value of $500,000 and remaining useful life of 5 years before an addition was made costing $50,000 that extended its useful life 3 more years. What is the new depreciation rate assuming the machine has no salvage value?
Answer #1

Book Value of Machine = $500,000

Add: Additional cost = $50,000

Total Cost (Revised) = $550,000

Estimated useful life (revised) = 5 + 3 = 8 years

New Depreciation rate = 1 / Estimated useful life

New Depreciation rate = 1 / 8

New Depreciation rate = 0.125

So, New Depreciation rate = 12.50%